# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ 2. Prerequisites
The purpose of this section is to develop (or supply) certain preliminaries which will be assembled in the bodies of proofs of main results.
@@ 2.1. Automorphic forms
for any $m,n\geq 1$ with $(n,M)=1$. In particular, $\psi_{f}(1)\lambda_{f}(n)=\psi(n)$, if $(n,M)=1$. Therefore,
$\overline{\lambda_{f}(n)}=\overline{\chi(n)}\lambda_{f}(n),\quad\lambda_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\lambda_{f}\left(\frac{mn}{d^{2}}\right),$ (2.9)
if $(mn,M)=1.$
The Hecke eigenbasis $\mathcal{B}_{k}(M,\chi)$ also contains a subset of newforms $\mathcal{B}^{\ast}_{k}(M,\chi)$, those forms which are simultaneous eigenfunctions of all the Hecke operators $T_{n}$ for any $n\geq 1$, and normalized to have first Fourier coefficient $\psi_{f}(1)=1$. The elements of $\mathcal{B}^{\ast}_{k}(M,\chi)$ are usually called primitive forms in the sense of Atkin-Lehner; that is, each one is orthogonal to all oldforms, and is an eigenfunction of all the Hecke and Atkin-Lehner operators.
@@ 2.2. Automorphic $L$-functions and functional equations
Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N)$ with trivial nebentypus, which has a Fourier-Whittaker expansion with the Fourier coefficients $A_{F}(m,n)$. The Fourier coefficients of $F$ and that of its contragredient $\widetilde{F}$ are related by $A_{F}(m,n)=A_{\widetilde{F}}(n,m)$ for any $(mn,N)=1$, with $A_{F}(1,1)=1$; see, for example, [11, Chapter 6] for definition and backgrounds. The Jacquet-Langlands $L$-function is given by $L(s,F)=\sum_{n\geq 1}A_{F}(n,1)n^{-s}$ for $\text{Re}(s)>1$. Now, let
@@
Next, for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we turn to considering the $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, which is given by the Dirichlet series $\sum_{m,n\geq 1}A_{F}(n,m)\lambda_{f}(n)(nm^{2})^{-s}$ for $\text{Re}(s)>1$. Now, one defines the Gamma factor $L_{\infty}(s,F\otimes f)$ by
$\begin{split}L_{\infty}(s,F\otimes f)&=\pi^{-3s}\prod_{i=1}^{3}\Gamma{\left(\frac{s+{(k-1)}/{2}+\alpha_{i}}{2}\right)}\Gamma{\left(\frac{s+{(k+1)}/{2}+\alpha_{i}}{2}\right)}\end{split}$
@@ 2.3. Epsilon constant
Let the parameters $k,N,M$ be as before. Define
@@
For any given holomorphic newform $f\in L^{2}(\Gamma_{0}(M)\backslash\text{SL}(2,\mathbb{R}))$ of weight $k$, let us define a function $\varphi_{f}\in L^{2}(\text{GL}(2,\mathbb{Q})\backslash\text{GL}(2,\mathbb{A}_{\mathbb{Q}})/\text{K}(2,M))$ by
@@
for $\gamma\in\text{GL}(2,\mathbb{Q})$, $g_{\infty}\in\text{GL}(2,\mathbb{R})$ and $k\in\text{K}(2,M)$, with $\text{K}(2,M)=\text{O}(2,\mathbb{R})\times\prod_{p|M}k_{2,p}\times\prod_{p\nmid M}\text{GL}(2,\mathbb{Z}_{p})$. It can be seen that each function $\varphi_{f}$ generates the automorphic representation $\pi_{f}$ of $\text{GL}(2,\mathbb{A}_{\mathbb{Q}})$ with trivial central character. We refer the reader to, for example, [12, Sections 13.3-13.6] for relevant backgrounds.
@@
for any $\gamma\in\text{GL}(3,\mathbb{Q})$, $g_{\infty}\in\text{GL}(3,\mathbb{R})$ and $k\in\text{K}(3,N)$, with $\text{K}(3,N)=\text{O}(3,\mathbb{R})\times\prod_{p|N}k_{3,p}$
×$\times\prod_{p\nmid N}\text{GL}(3,\mathbb{Z}_{p})$. It is well-known that one might factor
@@
We shall now develop some facts on the epsilon constants, which play a key rôle in later analysis. Observe that, for any unramified places $v=p$ with $p\nmid NM$, there already hold that $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=1$. We will focus on the archimedean place $v=\infty$ and the finite places $v=p,p|NM$ in the following.
@@
  • For $p|N$, one observes that the local component $\Pi_{F,p}$ is the Steinberg representation; while, $\pi_{g,p}$ is an unramified principal series representation with trivial central character, so that
@@ 2.4. Voronoĭ summation formula
$\Omega^{\pm}_{w}(x)=\frac{1}{2\pi i}\int_{(\sigma)}x^{-s}\gamma_{\pm}(s)\widetilde{w}(-s)\mathrm{d}s,$ (2.14)
@@
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{a\overline{n}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sqrt{N}\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am\overline{q},\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}Nm};w\right).\end{split}$ (2.15)
@@
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{a\overline{n}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am,\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}m};w\right).\end{split}$ (2.16)
One has a good control of the asymptotic behavior of $\Omega_{\pm}$, upon expressing it in term of the oscillatory integral; see, for example, [31, Lemma 6.1].
@@
$\frac{X^{{2}/{3}}}{cN^{1/6}}\sum_{\pm,n_{1}|c}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{(n_{2}/n_{1})^{1/{3}}}\,S(a,\pm n_{2};c/n_{1})\int_{\mathbb{R}^{+}}w(y)\,e\left(\pm\frac{3\left(n_{2}n^{2}_{1}Xy\right)^{1/{3}}}{c}\right)\mathrm{d}y.$ (2.17)
The main contribution occurs when $n_{2}\ll c^{3}/n^{2}_{1}X^{1-\varepsilon}$, by repeated integration by parts, and in which case $\Omega_{\pm}\ll X^{\varepsilon}$.
2.5. A Wilton-type bound
We record the following well-known Wilton-type bound for Hecke-Maaß forms on $\text{GL}(3)$; see [37, Theorem 1.1].
@@ 2.5. A Wilton-type bound
Let $X>0$. Let $W$ be a smooth function supported on $[1/2,5/2]$ with partial derivatives satisfying $x^{i}W(x/X)\ll_{i}ZP^{i}$ for any $i\geq 0$ and some $Z,P>0$. Then, for any $\alpha\in\mathbb{R}$ we have
$\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}e(m\alpha)\,W{\left(\frac{m}{X}\right)}\ll ZPX^{1/4+\varepsilon}.$
@@ 2.6. Petersson trace formula
where the spectral weights $\omega_{f}$ are given by $\omega_{f}:=\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\langle f,f\rangle$. In particular, if $f\in\mathcal{B}^{\ast}_{k}(M)$, one has $\omega_{f}=\frac{(k-1)M}{2\pi^{2}}L(1,\text{sym}^{2}f),$ which satisfies that $k^{1-\varepsilon}M\ll\omega_{f}\ll k^{1+\varepsilon}M$ for any $\varepsilon>0$; see [22, Lemma 2.5].
@@
For the purpose of our application, we need an expression for $\Delta^{\ast}_{k,M}$ instead of $\Delta_{k,M}$. Suppose that $m,n$ are positive integers satisfying $(m,M)=1$, we have the following identity (see [22, Eqn. (2.51)]):
@@ 2.7. Bounds on bilinear forms with Kloosterman sums
The bilinear form with Kloosterman sums always plays a vital rôle in many topics such as the moments of central values of $L$-functions, subconvexity even the classic bilinear exponential sums; see, for example, [28, 24, 44] for descriptions in different regards. In this paper, we will develop the following relevant results to facilitate our use after a while.
@@
Let $X,Y\geq 1$. Let $W$ be a smooth function supported on $[1/2,5/2]\times[1/2,5/2]$ with partial derivatives satisfying $x^{i}y^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial y^{j}}W(x/X,y/Y)\ll_{i,j}ZZ^{i}_{1}Z^{j}_{2}$ for some $Z,Z_{1},Z_{2}>0$. Let $q\geq 1$ be an integer. For any $h\in\mathbb{Z}$, with $(h,q)=1$, we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)\,W{\left(\frac{n}{X},\frac{m}{Y},\right)}\ll Z_{1}Z\sqrt{XYq}+Z_{1}ZXq^{3/4}+Z_{1}Zq.$ (2.18)
@@
This lemma essentially follows from the work of Kerr, Shparlinski, Wu and Xi [24] which appealed to the approach in [44, Theorem 2.1]. By partial summation, one finds that
$\begin{split}&\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\\ &\ll\int^{5X/2}_{X/2}\left|\hskip 5.0pt\sideset{}{{}^{\ast}}{\sum}\limits_{x\bmod q}\sum_{m\geq 1}\alpha_{m,t}\,e{\left(\frac{m\overline{x}}{q}\right)}\sum_{n\leq t}e{\left(\frac{nhx}{q}\right)}\right|\mathrm{d}t,\end{split}$
with $\alpha_{m,t}=\frac{\partial}{\partial t}W(t/X,m/Y)A_{F}(m,1)/\sqrt{m}$. This is compatible with the setting as presented in the beginning of the [44, Section 5]. We still might bound the inner triple sum by ±$\sum_{\pm}\sum_{1\leq i\leq I}|S^{\pm}_{i}|$, where, for $1\leq i\leq I$, $S^{\pm}_{i}=\sum_{m\leq X}\alpha_{m,t}\gamma_{x}\,e{({\pm m\overline{x}}/{q})}$; see [24, Section 6] for details. ∎
Observe that the bilinear forms we are concerned about is a smooth-weighted sum. The direct use of the Possion shows some bonus. Indeed, we alternatively have
@@
Keep the notation as in Lemma 2.6. Then we have
@@
$\begin{split}&\ll\left(\frac{Y}{q}\sum_{n_{1},n_{2}\geq 1}\,\sum_{|\jmath|\ll Z_{2}q/Y^{1-\varepsilon}}\Phi(\jmath;n_{1},n_{2})\sum_{\alpha\bmod q}S(n_{1},\alpha;q)\overline{S(n_{2},\alpha;q)}\,e{\left(\frac{\jmath\alpha}{q}\right)}\right)^{1/2},\end{split}$
with $\Phi(\jmath;v_{1},v_{2})=\int_{\mathbb{R}^{+}}W(v_{1},t)W(v_{2},t)e(-\jmath t/q)\mathrm{d}t$ for any $v_{1},v_{2}\in\mathbb{R}^{+}$ and $\jmath\in\mathbb{R}$. It is verifiable that the zero-frequence shows a quantity $Z\sqrt{XYq}+XZ\sqrt{Y}$; while, if $Z_{2}q>Y$, the non-zero frequences, however, will contribute the quantity $\sqrt{Z_{2}}Zq^{3/4}\sqrt{Y}$ by invoking the estimate for triple sums of Kloosterman sums; see, for example, [24, Lemma 4.1]. ∎

### v4

@@ 2. Prerequisites
The purpose of this section is to develop (or supply) certain preliminaries which will be assembled in the bodies of proofs of the main results.
@@ 2.1. Automorphic forms
for any $m,n\geq 1$ with $(n,M)=1$. In particular, $\psi_{f}(1)\lambda_{f}(n)=\psi(n)$, if $(n,M)=1$. Therefore, if $(mn,M)=1$,
$\overline{\lambda_{f}(n)}=\overline{\chi(n)}\lambda_{f}(n),\quad\lambda_{f}(m)\lambda_{f}(n)=\sum_{d|(n,m)}\chi(d)\lambda_{f}\left(\frac{mn}{d^{2}}\right).$ (2.2)
The Hecke eigenbasis $\mathcal{B}_{k}(M,\chi)$ also contains a subset of newforms $\mathcal{B}^{\ast}_{k}(M,\chi)$, those forms which are simultaneous eigenfunctions of all the Hecke operators $T_{n}$ for any $n\geq 1$, and normalized to have first Fourier coefficient $\psi_{f}(1)=1$. The elements of $\mathcal{B}^{\ast}_{k}(M,\chi)$ are usually called primitive forms in the sense of Atkin-Lehner; that is, each one is orthogonal to all oldforms, and is an eigenfunction of all the Hecke and Atkin-Lehner operators. For convenience, we abbreviate the notation $\mathcal{B}^{\ast}_{k}(M)$ to $\mathcal{B}^{\ast}_{k}$, the space of newforms of weight $k$ and arbitrary level parameter, if one does not specify the level parameter.
@@ 2.2. Automorphic $L$-functions and functional equations
Let $F(z)$ be a normalized Hecke-Maaß form of type $\nu=(\nu_{1},\nu_{2})$ for the congruent subgroup $\Gamma_{0}(N)$ with trivial nebentypus, which has a Fourier-Whittaker expansion with the Fourier coefficients $A_{F}(m,n)$. The Fourier coefficients of $F$ and that of its contragredient $\widetilde{F}$ are related by $A_{F}(m,n)=A_{\widetilde{F}}(n,m)$ for any $(mn,N)=1$, with $A_{F}(1,1)=1$; see, for example, [13, Chapter 6] for definition and backgrounds. The Jacquet-Langlands $L$-function is given by $L(s,F)=\sum_{n\geq 1}A_{F}(n,1)n^{-s}$ for $\text{Re}(s)>1$. Now, let
@@
Next, for any $f\in\mathcal{B}^{\ast}_{k}(M)$, we turn to considering the $\text{GL}(3)\times\text{GL}(2)$ Rankin-Selberg $L$-function $L(s,F\otimes f)$, which is given by the Dirichlet series
$L(s,F\otimes f)=\sum_{m,n\geq 1}\frac{A_{F}(n,m)\lambda_{f}(n)}{(nm^{2})^{s}}$
for $\text{Re}(s)>1$. Now, one defines the Gamma factor $L_{\infty}(s,F\otimes f)$ by
$\begin{split}L_{\infty}(s,F\otimes f)&=\pi^{-3s}\prod_{i=1}^{3}\Gamma{\left(\frac{s+{(k-1)}/{2}+\alpha_{i}}{2}\right)}\Gamma{\left(\frac{s+{(k+1)}/{2}+\alpha_{i}}{2}\right)}\end{split}$ (2.5)
@@ 2.3. Epsilon constant
In this part, we will be dedicated to developing some facts on the epsilon constant $\varepsilon(F\otimes f)$, which play a key rôle in later analysis. We proceed by borrowing some points of view from the automorphic representations of algebraic groups over local fields. Let the parameters $k,N,M$ be as before. Define
@@
For any given holomorphic newform $f\in L^{2}(\Gamma_{0}(M)\backslash\text{SL}(2,\mathbb{R}))$ of weight $k$, let us proceed by defining a function $\varphi_{f}\in L^{2}(\text{GL}(2,\mathbb{Q})\backslash\text{GL}(2,\mathbb{A}_{\mathbb{Q}})/\text{K}(2,M))$ by
@@
for $\gamma\in\text{GL}(2,\mathbb{Q})$, $g_{\infty}\in\text{GL}(2,\mathbb{R})$ and $k\in\text{K}(2,M)$, with
$\text{K}(2,M)=\text{O}(2,\mathbb{R})\times\prod_{p|M}k_{2,p}\times\prod_{p\nmid M}\text{GL}(2,\mathbb{Z}_{p}).$
It can be seen that each function $\varphi_{f}$ generates the automorphic representation $\pi_{f}$ of $\text{GL}(2,\mathbb{A}_{\mathbb{Q}})$ with trivial central character. We refer the reader to, for example, [14, Sections 13.3-13.6] for relevant backgrounds.
@@
for any $\gamma\in\text{GL}(3,\mathbb{Q})$, $g_{\infty}\in\text{GL}(3,\mathbb{R})$ and $k\in\text{K}(3,N)$, with
$\text{K}(3,N)=\text{O}(3,\mathbb{R})\times\prod_{p|N}k_{3,p}\times\prod_{p\nmid N}\text{GL}(3,\mathbb{Z}_{p}).$
It is well-known that one might factor
@@
We shall now ready to deduce an explicit form of the epsilon constant $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})$ at each place $p\leq\infty$. Observe that, for any unramified places $v=p$ with $p\nmid NM$, there already hold that $\varepsilon(\Pi_{F,p}\otimes\pi_{f,p})=1$. We will focus on the archimedean place $v=\infty$ and the finite places $v=p,p|NM$ in the following.
@@
  • For $p|N$, one observes that the local component $\Pi_{F,p}$ is the Steinberg representation; whilst, $\pi_{g,p}$ is an unramified principal series representation with trivial central character, so that
@@ 2.4. Voronoĭ summation formula
$\Omega_{\pm}(x;w)=\frac{1}{2\pi i}\int_{(\sigma)}x^{-s}\gamma_{\pm}(s)\widetilde{w}(-s)\mathrm{d}s,$ (2.8)
@@
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{n\overline{a}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sqrt{N}\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am\overline{N},\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}Nm};w\right).\end{split}$ (2.9)
@@
    $\begin{split}&\sum_{n>0}A_{F}(m,n)e\left(\frac{n\overline{a}}{c}\right)w\left(\frac{n}{X}\right)\\ &\hskip 34.14322pt=c\sum_{\pm}\sum_{n_{1}|cm}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{n_{1}n_{2}}S(am,\pm n_{2};cm/n_{1})\,\Omega_{\pm}\left(\frac{n_{2}n^{2}_{1}X}{c^{3}m};w\right).\end{split}$ (2.10)
One has a good control of the asymptotic behavior of $\Omega_{\pm}$, upon expressing it in term of the oscillatory integral; see, for example, [34, Lemma 6.1].
@@
$\frac{X^{{2}/{3}}}{cN^{1/6}m^{2/3}}\sum_{\pm,n_{1}|c}\sum_{n_{2}>0}\frac{\overline{A_{F}(n_{1},n_{2})}}{(n_{2}/n_{1})^{1/{3}}}\,S(am\overline{N},\pm n_{2};cm/n_{1})$
$\hskip 179.25244pt\int_{\mathbb{R}^{+}}w(y)\,e\left(\pm\frac{3\left(n_{2}n^{2}_{1}Xy/m\right)^{1/{3}}}{c}\right)\mathrm{d}y.$ (2.11)
The main contribution occurs when $n_{2}\ll c^{3}m/n^{2}_{1}X^{1-\varepsilon}$, by repeated integration by parts, and in which case $\Omega_{\pm}\ll X^{\varepsilon}$.
2.5. Non-linear exponential sums
We record the following square-root cancellation estimate for the non-linear exponential sums related to Hecke-Maaß forms on $\text{GL}(3)$; see [45, Theorem 3] and the remarks after [45, Eqn. (3.18)].
@@ 2.5. Non-linear exponential sums
Let $X\geq 2$. Let $W$ be a smooth function supported on $[1/2,5/2]$ with partial derivatives satisfying $x^{i}W(x/X)\ll_{i}ZP^{i}$ for any $i\geq 0$ and some $Z,P>0$. Then, for any $\alpha\in\mathbb{R}$ with $\alpha^{2}\in\mathbb{Q}$, we have
$\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}e(\alpha\sqrt{m})\,W{\left(\frac{m}{X}\right)}\ll ZP(1+|\alpha|\log X).$
@@ 2.6. Petersson trace formula
where the spectral weights $\omega_{f}$ are given by $\omega_{f}:=\frac{(4\pi)^{k-1}}{\Gamma(k-1)}\langle f,f\rangle$. In particular, if $f\in\mathcal{B}^{\ast}_{k}(M)$, one has $\omega_{f}=\frac{(k-1)M}{2\pi^{2}}L(1,\text{sym}^{2}f),$ which satisfies that $k^{1-\varepsilon}M\ll\omega_{f}\ll k^{1+\varepsilon}M$ for any $\varepsilon>0$; see [24, Lemma 2.5].
@@
For the purpose of our application, we need an expression for $\Delta^{\ast}_{k,M}$ instead of $\Delta_{k,M}$. Suppose that $m,n$ are positive integers satisfying $(m,M)=1$, we have the following identity (see [24, Eqn. (2.51)]):
@@ 2.7. Bounds on bilinear forms with Kloosterman sums
The bilinear form with Kloosterman sums always plays a vital rôle in many topics such as the moments of central values of $L$-functions, subconvexity even the classical bilinear exponential sums in additive number theory; see, for example, [31, 26, 47] for descriptions in different regards. In the present paper, we will develop the following relevant results to facilitate our use after a while.
@@
Let $X,Y\geq 1$. Let $W$ be a smooth function supported on $[1/2,5/2]\times[1/2,5/2]$ with partial derivatives satisfying $x^{i}y^{j}\frac{\partial^{i}}{\partial x^{i}}\frac{\partial^{j}}{\partial y^{j}}W(x/X,y/Y)\ll_{i,j}ZZ^{i}_{1}Z^{j}_{2}$ for some $Z,Z_{1},Z_{2}>0$. Let $q\geq 1$ be an integer. For any $h\in\mathbb{Z}$, we have
$\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(nh,m;q)\,W{\left(\frac{n}{X},\frac{m}{Y},\right)}\ll Z_{1}Z\sqrt{XYq}+Z_{1}ZXq^{3/4}(h,q)^{1/4}+Z_{1}Zq.$ (2.12)
@@
This lemma essentially follows from the work of Kerr, Shparlinski, Wu and Xi [26] which appealed to the approach in [47, Theorem 2.1]. By partial summation, one finds that
$\begin{split}&\sum_{n\geq 1}\sum_{m\geq 1}\frac{A_{F}(m,1)}{\sqrt{m}}S(n,hm;q)W{\left(\frac{n}{X},\frac{m}{Y}\right)}\\ &\hskip 85.35826pt\ll\int^{5X/2}_{X/2}\left|\hskip 5.0pt\sideset{}{{}^{\ast}}{\sum}\limits_{x\bmod q}\sum_{m\geq 1}\alpha_{m,t}\,e{\left(\frac{m\overline{x}}{q}\right)}\sum_{n\leq t}e{\left(\frac{nhx}{q}\right)}\right|\mathrm{d}t,\end{split}$
with $\alpha_{m,t}=\frac{\partial}{\partial t}W(t/X,m/Y)A_{F}(m,1)/\sqrt{m}$. This is compatible with the setting as presented in the beginning of the [47, Section 5]. We still might bound the inner triple sum by ±$\sum_{\pm}\sum_{1\leq i\leq I}|S^{\pm}_{i}|$, where, for $1\leq i\leq I$, $S^{\pm}_{i}=\sum_{m\leq X}\alpha_{m,t}\gamma_{x}\,e{({\pm m\overline{x}}/{q})}$; see [26, Section 6] for details. ∎
Observe that the bilinear forms we are concerned about is a smooth-weighted sum. A direct use of the Possion might show some bonus. Indeed, we alternatively have
@@
Keep the notation as in Lemma 2.6. Then, for any $(h,q)=1$, we have
@@
$\begin{split}&\ll\left(\frac{Y}{q}\sum_{n_{1},n_{2}\geq 1}\,\sum_{|\jmath|\ll Z_{2}q/Y^{1-\varepsilon}}\Phi(\jmath;n_{1},n_{2})\sum_{\alpha\bmod q}S(n_{1}h,\alpha;q)\overline{S(n_{2}h,\alpha;q)}\,e{\left(\frac{\jmath\alpha}{q}\right)}\right)^{1/2},\end{split}$
with $\Phi(\jmath;v_{1},v_{2})=\int_{\mathbb{R}^{+}}W(v_{1},t)W(v_{2},t)e(-\jmath t/q)\mathrm{d}t$ for any $v_{1},v_{2}\in\mathbb{R}^{+}$ and $\jmath\in\mathbb{R}$. It is verifiable that the zero-frequence shows a quantity $Z\sqrt{XYq}+XZ\sqrt{Y}$; while, if $Z_{2}q>Y$, the non-zero frequences, however, will contribute the quantity $\sqrt{Z_{2}}Zq^{3/4}\sqrt{Y}$ by invoking the estimate for triple sums of Kloosterman sums; see, for example, [26, Lemma 4.1]. ∎
It is remarkable that, as it can be seen in (2.13), the upper-bound becomes smaller, as $m,n$ vary in the ranges with relatively small heights.